# Goal-Reaching Trajectory Design Near Danger with Piecewise Affine Reach-avoid Computation

### v3

@@ III-C Reach-Avoid Problem
Note, we only consider static obstacles, because our goal is to plan a robot’s motion near danger. The danger resulting from moving or adversarial obstacles depends upon accurately predicting obstacle motion, which introduces an independent variable unrelated to our purpose. We leave extension to dynamic scenes for future work, but we anticipate we can adapt reachability-based methods such as [vaskov2019towards, vaskov2019not].
@@
$\begin{split}\mathrm{\textnormal{BRAS}}(t_{\mathrm{\textnormal{f}}},\overline{P_{\mathrm{\textnormal{G}}}},\overline{\mathcal{O}})=\{&(\mathbf{p}_{0},\mathbf{k})\in P\times K\ \mid\forall~{}\mathbf{z}_{0}~{}\mathrm{\textnormal{s.t.}}\\ &(\mathbf{z}_{0})_{1:{n}_{w}}=\mathbf{p}_{0},\\ &(\mathbf{z}(t_{\mathrm{\textnormal{f}}};\mathbf{z}_{0},\mathbf{k}))_{1:{n}_{w}}\in\overline{P_{\mathrm{\textnormal{G}}}},~{}\mathrm{\textnormal{and}}\\ &(\mathbf{z}(\tau;\mathbf{z}_{0},\mathbf{k}))_{1:{n}_{w}}\notin\overline{\mathcal{O}}\quad\forall\tau\in[0,t_{\mathrm{\textnormal{f}}}]\}\end{split}$ (15)
@@
We restrict the goal and obstacles to the workspace only to simplify exposition; we demonstrate more general goals and obstacles, such as in the robot’s orientation, in Sections V and VI.
@@
To simplify exposition, we treat tracking error as a lumped uncertainty representation between planning and control, encompassing multiple error sources – disturbances, state estimation, object detection, and trajectory prediction. Separately analyzing these sources would introduce confounding variables, complicating the interpretation of our results. Instead, we consolidate them into a single metric of tracking error. For example, one can design a robust tracking controller that yields a tracking error bound, robust to external disturbances and model uncertainties [chen2021fastrack, bechlioulis2014low]. Our planning level only considers the result of those errors as a tracking error bound without separately accounting for each. Alternatively, if the bounds of inertial uncertainties [michaux2023can], dynamic disturbance [selim2022safe, alanwar2023data], and perception uncertainty [kousik2020bridging] are known (which are strong assumptions), then planner-tracker frameworks similar to ours have been shown to maintain reach-avoid guarantees in the past by “buffering” the bounds into the tracking error [kousik2020bridging, vaskov2019towards, selim2022safe, chen2021fastrack]. In Section VII, we show the preliminary results of a hardware experiment that successfully includes localization drift from the inertial measurement unit (IMU) and the encoders.
@@ III-D Example System: TurtleBot
with workspace (position) states $\mathbf{w}=[p_{x},p_{y}]^{\intercal}$, other (heading) state $\mathbf{p}_{\mathrm{\textnormal{other}}}=\theta$, and trajectory parameters $\mathbf{k}=[v_{\mathrm{\textnormal{des}}},\omega_{\mathrm{\textnormal{des}}}]^{\intercal}$. Note that a Dubins car model (16) is ETI in the workspace (Assum. 2) since its dynamics in workspace dimension is translation-invariant (i.e. $\left\|\tfrac{\partial\mathbf{f}_{\mathbf{w}}}{\partial\mathbf{w}}\right\|_{2}=0$.)
@@
where $\mathbf{z}_{\mathrm{\textnormal{other}}}=[\theta,v]^{\intercal}$ are heading and linear velocity, and $\mathbf{u}_{\mathrm{\textnormal{fb}}}=[u_{\omega},u_{v}]^{\intercal}$ is implemented as an iterative LQR tracking controller [kong2021ilqr] that commands angular velocity and linear acceleration.
We define a goal set as a unit box, $\overline{P_{\mathrm{\textnormal{G}}}}=\{\mathbf{w}\mid p_{x},p_{y}\in[-1,1]\}$, and an obstacle near the goal, $\overline{\mathcal{O}}=\{\mathbf{w}\mid p_{x}\in[-1.75,-1.25],p_{y}\in[-0.25,0.25]\}$ (see Figures 2 and 3).
@@ IV-A1 Augmented State, Goal, and Obstacles
Before proceeding, to enable PARC, we define an augmented planning state $\mathbf{x}(t)\in X=W\times K\times P_{\mathrm{\textnormal{other}}}\subset\mathbb{R}^{{n}_{w}+{n}_{k}+{{n}_{p}}_{\mathrm{\textnormal{other}}}}$ at time $t$ as:
@@
We perform this reordering and stacking because, as we will see later in Lem. 12, the $\mathbf{w}$ and $\mathbf{k}$ are ETI states, so indexing them as the first states of $\mathbf{x}$ will simplify avoid set computation. We similarly augment the goal set $\overline{P_{\mathrm{\textnormal{G}}}}$ and the obstacles $\overline{\mathcal{O}}=\bigcup_{i=1}^{{n}_{\mathcal{O}}}\overline{\mathcal{O}}_{i}$, which we assume are union of H-polytopes:
$\begin{split}P_{\mathrm{\textnormal{G}}}&=\overline{P_{\mathrm{\textnormal{G}}}}\times K\times P_{\mathrm{\textnormal{other}}}\\ \mathcal{O}_{i}&=\overline{\mathcal{O}}_{i}\times K\times P_{\mathrm{\textnormal{other}}}.\end{split}$ (19)
@@ IV-A2 PWA Planning Model
We approximate the planning model $\mathbf{f}_{\mathrm{\textnormal{plan}}}$ as a discrete-time, time-variant PWA to enable computation. First, we define linearization points $\mathbf{x}^{*}_{t,i}\in X_{t}$, $i=1,\cdots,{n}^{*}_{t}$ for each $t=0,\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t$, which can be generated by gridding or uniform sampling. We define the PWA regions as the Voronoi partition of the linearization points [casselman2009new] (see Appendix -C). Each $i$th region is an H-polytope $\mathcal{P}(\mathbf{A}_{i,t},\mathbf{b}_{i,t})$. Then, in each region, at each time $t=0,\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t$, we define the dynamics $\mathbf{C}_{i,t}$ and $\mathbf{d}_{i,t}$ by Taylor expanding about the linearization points:
$\mathbf{x}(t+\Delta t)=\mathbf{g}_{\mathrm{\textnormal{plan}},t}(t,\mathbf{x}^{*}_{i,t})+\frac{\partial\mathbf{g}_{\mathrm{\textnormal{plan}},t}(t,\mathbf{x}(t))}{\partial\mathbf{x}(t)}\Bigr{|}_{\mathbf{x}={\mathbf{x}^{*}_{i,t}}}(\mathbf{x}(t)-\mathbf{x}^{*}_{i,t}),$ (20a)
$\mathbf{g}_{\mathrm{\textnormal{plan}},t}(t,\mathbf{x}(t))=\begin{bmatrix}(\mathbf{p}(t)+\Delta t\mathbf{f}_{\mathrm{\textnormal{plan}}}(t,\mathbf{p}(t),\mathbf{k}))_{1:{n}_{w}}\\ \mathbf{k}\\ (\mathbf{p}(t)+\Delta t\mathbf{f}_{\mathrm{\textnormal{plan}}}(t,\mathbf{p}(t),\mathbf{k}))_{{n}_{w}+1:{n}_{p}}\\ \end{bmatrix}$ (20b)
$\mathbf{C}_{i,t}=\frac{\partial\mathbf{g}_{\mathrm{\textnormal{plan}},t}(t,\mathbf{x}(t))}{\partial\mathbf{x}(t)}\Bigr{|}_{{\mathbf{x}=\mathbf{x}^{*}_{i,t}}},\quad\mathrm{\textnormal{and}}$ (20c)
$\mathbf{d}_{i,t}=\mathbf{g}_{\mathrm{\textnormal{plan}},t}(t,\mathbf{x}^{*}_{i,t})-\frac{\partial\mathbf{g}_{\mathrm{\textnormal{plan}},t}(t,\mathbf{x}(t))}{\partial\mathbf{x}(t)}\Bigr{|}_{{\mathbf{x}=\mathbf{x}^{*}_{i,t}}}\mathbf{x}^{*}_{i,t}.$ (20d)
@@
Choosing a larger $\Delta t$ would induce higher mismatch between the tracking model and the planning model, thus increasing the conservativeness of the method. In scenarios with highly dynamic systems or cluttered environments, selecting a smaller $\Delta t$ reduces conservatism, as demonstrated in Fig. 4. As will be shown in the rest of this section, PARC’s reach-avoid guarantees will still be maintained regardless of the choice of $\Delta t$, and its computation time is inversely proportional to $\Delta t$.
@@ IV-B PARC: Reach Set without Tracking Error
Without accounting for tracking error, the reach set of a planning model is the set of initial conditions $\mathbf{x}_{0}$ of trajectories that are guaranteed to reach the goal region $P_{\mathrm{\textnormal{G}}}$ at time $t_{\mathrm{\textnormal{f}}}$. Instead of naively computing the full reach set, which scales exponentially in memory, computation time, and a number of set representations with the number of timesteps, PARC is planning aware: given an expert plan, we compute only a local, underapproximated BRS efficiently for the neighborhood that shares the same mode sequence as the expert plan. Fixing the mode sequence reduces the PWA system to a time-variant affine system, so the BRS is a single H-polytope using the inverse affine map of the goal (see (9)), with computation time scaling linearly with the number of timesteps and quadratically with dimensions from matrix multiplication.
@@
$\mathbf{\psi}\!\left(\mathbf{x}_{0}\right)\in P_{\mathrm{\textnormal{G}}}.$ (21)
Obtaining expert plans is system-dependent, but is not difficult, because we do not require the expert plan to avoid obstacles. Thus, one option is to solve a two-point boundary value nonlinear MPC problem, for which solution guarantees exist [bonalli2019gusto]. In practice, it is often sufficient to uniformly sample trajectory parameters $\mathbf{k}$ and roll out the PWA planning model, since the robot starts out near the goal. This assumption is even easier to fulfill if the planning model is time-variant affine (which we use for a general 3D quadrotor and drift vehicle planning in Section V, VI) since plans have only one mode.
@@
$s_{t}$ $=\min\{i\ |\ \mathbf{x}(t)\in\mathcal{P}(\mathbf{A}_{i,t},\mathbf{b}_{i,t}),i=1,\dots,{n}_{\mathrm{\textnormal{PWA}},t}\},$ (22)
@@
$\begin{split}\overline{\Omega}_{t_{\mathrm{\textnormal{f}}}}&=P_{\mathrm{\textnormal{G}}},\\ \overline{\Omega}_{t}&=\mathcal{B}\!\left(\overline{\Omega}_{t+\Delta t},\mathbf{C}_{s_{t},t},\mathbf{d}_{s_{t},t}\right)\end{split}$ (23)
@@
This follows directly from Lemma 1. ∎
@@
Traditionally, methods such as the reachableSet function in MPT3 [herceg2013multi] compute the full BRS of the PWA system instead of that of only a particular mode sequence. While computing the full BRS would return all possible reachable states and do not require knowledge of an expert plan, the computation time grows exponentially with the number of BRS steps and quickly becomes intractable due to the large number of polytopes, as shown in Table I. Instead, Proposition 10 boasts a much more reasonable computation time by isolating the affine dynamics of a single mode sequence and keeping only one polytope regardless of the number of BRS steps.
Figure 2: (Right) Visualization of the reach set computation in Section IV-B for the TurtleBot example (Section III-D), projected onto the planning state space $P$, when $\mathbf{p}_{0}=[-4,0,\pi/5]^{\intercal}$, $t_{\mathrm{\textnormal{f}}}=4(s)$. The goal set is the green polytope extended to all $\theta\in[-\pi,\pi]$. An expert plan, computed by solving a two-point boundary value problem, passes through two modes (blue and magenta line). We compute the corresponding $t_{\mathrm{\textnormal{f}}}$-time BRS $\overline{\Omega}_{0}$ (blue polytope with solid outline). Intermediate BRSs $\overline{\Omega}_{t}$ are the polytopes without outlines, colored by mode. (Left) Illustration of the trajectory parameter space; the grey region represents $K$, the green region represents the projection of $\overline{\Omega}_{0}$ on $K$. After reach-set computation, the state $\mathbf{x}_{\mathrm{\textnormal{sample}}}=[\mathbf{p}_{\mathrm{\textnormal{sample}}}^{\intercal},\mathbf{k}_{\mathrm{\textnormal{sample}}}^{\intercal}]^{\intercal}$ is sampled from $\overline{\Omega}_{0}$: black dot (right) denotes the $\mathbf{p}_{\mathrm{\textnormal{sample}}}$ and black star (left) denotes the $\mathbf{k}_{\mathrm{\textnormal{sample}}}$. The resulting trajectory is denoted as the black line which reaches the goal.
@@
TABLE I: Computation time and number of polytopes for 1, 2, 3, 4, 5, and 26-step BRS for the TurtleBot example (Section III-D) using the reachableSet function in MPT3 [herceg2013multi] and Proposition 10. Since MPT3 computes the full BRS of the PWA system, the computation time quickly grows intractable due to the large number of polytopes. In contrast, Proposition 10 maintains a reasonable computation time by keeping only one polytope at each step, characterized by the affine dynamics corresponding to the mode sequence of the expert plan.

### v4

@@ III-C Reach-Avoid Problem
Note, we only consider static obstacles, because our goal is to plan a robot’s motion near danger. The danger resulting from moving or adversarial obstacles depends upon accurately predicting obstacle motion, which introduces an independent variable unrelated to our purpose. We leave an extension to dynamic scenes for future work, but we anticipate that it can be accomplished by adapting reachability-based methods such as [vaskov2019towards, vaskov2019not].
@@
$\begin{split}\mathrm{\textnormal{BRAS}}(t_{\mathrm{\textnormal{f}}},\overline{P_{\mathrm{\textnormal{G}}}},\overline{\mathcal{O}})=\{&(\mathbf{p}_{0},\mathbf{k})\in P\times K\ \mid\forall~{}\mathbf{z}_{0}~{}\mathrm{\textnormal{s.t.}}\\ &(\mathbf{z}_{0})_{1:{n}_{p}}=\mathbf{p}_{0},\\ &(\mathbf{z}(t_{\mathrm{\textnormal{f}}};\mathbf{z}_{0},\mathbf{k}))_{1:{n}_{w}}\in\overline{P_{\mathrm{\textnormal{G}}}},~{}\mathrm{\textnormal{and}}\\ &(\mathbf{z}(\tau;\mathbf{z}_{0},\mathbf{k}))_{1:{n}_{w}}\notin\overline{\mathcal{O}}\quad\forall\tau\in[0,t_{\mathrm{\textnormal{f}}}]\}\end{split}$ (15)
@@
We restrict the goal and obstacles to the workspace only to simplify exposition; we demonstrate more general goals and obstacles, such as in the robot’s orientation, in Sections VI.
@@
To simplify exposition, we treat tracking error as a lumped uncertainty representation between planning and control, encompassing multiple error sources—disturbances, state estimation, object detection, and trajectory prediction. Separately analyzing these sources would introduce confounding variables, complicating the interpretation of our results. Instead, we consolidate them into a single metric of tracking error. For example, one can design a robust tracking controller that yields a tracking error bound, robust to external disturbances and model uncertainties [chen2021fastrack, bechlioulis2014low]. Our planning level only considers the result of those errors as a tracking error bound without separately accounting for each. Alternatively, if the bounds of inertial uncertainties [michaux2023can], dynamic disturbance [selim2022safe, alanwar2023data], and perception uncertainty [kousik2020bridging] are known (which are strong assumptions), then planner-tracker frameworks similar to ours have been shown to maintain reach-avoid guarantees in the past by “buffering” the bounds into the tracking error [kousik2020bridging, vaskov2019towards, selim2022safe, chen2021fastrack]. In Section VII, we show the preliminary results of a hardware experiment that successfully includes localization drift from the inertial measurement unit (IMU) and the encoders.
@@ III-D Example System: TurtleBot
with workspace (position) states $\mathbf{w}=[p_{x},p_{y}]^{\intercal}$, other (heading) state $\mathbf{p}_{\mathrm{\textnormal{other}}}=\theta$, and trajectory parameters $\mathbf{k}=[v_{\mathrm{\textnormal{des}}},\omega_{\mathrm{\textnormal{des}}}]^{\intercal}$ representing desired linear velocity and angular velocity. Note that a Dubins car model (16) is ETI in the workspace (Assum. 2) since changing the workspace states does not affect the planning dynamics (i.e. $\left\|\tfrac{\partial\mathbf{f}_{\mathbf{w}}}{\partial\mathbf{w}}\mathbf{f}_{\mathbf{w}}\right\|_{2}=\left\|\tfrac{\partial\mathbf{f}_{\mathbf{w}}}{\partial\mathbf{w}}\tfrac{\partial\mathbf{f}_{\mathbf{w}}}{\partial\mathbf{p}}\right\|_{\mathrm{\textnormal{F}}}=\left\|\tfrac{\partial\mathbf{f}_{\mathbf{w}}}{\partial\mathbf{w}}\tfrac{\partial\mathbf{f}_{\mathbf{w}}}{\partial\mathbf{k}}\right\|_{\mathrm{\textnormal{F}}}=0$.)
@@
where $\mathbf{z}_{\mathrm{\textnormal{other}}}=[\theta,v]^{\intercal}$ are heading and linear velocity, and $\mathbf{u}_{\mathrm{\textnormal{fb}}}=[u_{\omega},u_{v}]^{\intercal}$ is implemented as an iterative linear quadratic regulator (iLQR) tracking controller [kong2021ilqr] that commands angular velocity and linear acceleration.
We define a goal set as a unit box, $\overline{P_{\mathrm{\textnormal{G}}}}=\{\mathbf{w}\mid p_{x},p_{y}\in[-1,1]\}$, and an obstacle near the goal, $\overline{\mathcal{O}}=\{\mathbf{w}\mid p_{x}\in[-1.75,-1.25],p_{y}\in[-0.25,0.25]\}$ (see Fig. 2 and 3).
@@ IV-A1 Augmented State, Goal, and Obstacles
Before proceeding, to enable PARC, we define an augmented planning state $\mathbf{x}(t)\in X=W\times K\times P_{\mathrm{\textnormal{other}}}\subset\mathbb{R}^{{n}_{w}+{n}_{k}+{{n}_{p}}_{\mathrm{\textnormal{other}}}}=\mathbb{R}^{{n}_{x}}$ at time $t$ as:
@@
We perform this reordering and stacking because, as we will see later in Lem. 11, $\mathbf{w}$ and $\mathbf{k}$ are ETI states, so indexing them as the first states of $\mathbf{x}$ will simplify avoid set computation. We similarly augment the goal set $\overline{P_{\mathrm{\textnormal{G}}}}$ and the obstacles $\overline{\mathcal{O}}=\bigcup_{i=1}^{{n}_{\mathcal{O}}}\overline{\mathcal{O}}_{i}$, which we assume are union of H-polytopes:
$\begin{split}P_{\mathrm{\textnormal{G}}}&=\overline{P_{\mathrm{\textnormal{G}}}}\times K\times P_{\mathrm{\textnormal{other}}}\\ \mathcal{O}_{i}&=\overline{\mathcal{O}}_{i}\times K\times P_{\mathrm{\textnormal{other}}},\end{split}$ (19)
where $i=1,\cdots,{n}_{\mathcal{O}}$.
@@ IV-A2 PWA Planning Model
We approximate the planning model $\mathbf{f}_{\mathrm{\textnormal{plan}}}$ as a discrete-time, time-variant PWA to enable computation. First, we define linearization points $\mathbf{x}^{*}_{t,i}\in X$, $i=1,\cdots,{n}^{*}_{t}$ for each $t=0,\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t$, which can be generated by gridding or uniform sampling. We define the PWA regions as the Voronoi partition of the linearization points [casselman2009new] (see Appendix -C). Each $i$th region is an H-polytope $\mathcal{P}(\mathbf{A}_{i,t},\mathbf{b}_{i,t})$. Then, in each region, at each time $t=0,\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t$, we define the dynamics $\mathbf{C}_{i,t}$ and $\mathbf{d}_{i,t}$ by Taylor expanding about the linearization points:
$\mathbf{x}(t+\Delta t)$ $=\mathbf{C}_{i,t}\mathbf{x}(t)+\mathbf{d}_{i,t},$ (20a)
$\mathbf{C}_{i,t}$ $=\mathbf{I}_{{n}_{x}}+\Delta t\frac{\partial\mathbf{g}_{\mathrm{\textnormal{plan}}}(t,\mathbf{x})}{\partial\mathbf{x}}\Bigr{|}_{{\mathbf{x}=\mathbf{x}^{*}_{i,t}}},$ (20b)
$\mathbf{d}_{i,t}$ $=\Delta t(\mathbf{g}_{\mathrm{\textnormal{plan}}}(t,\mathbf{x}^{*}_{i,t})-\frac{\partial\mathbf{g}_{\mathrm{\textnormal{plan}}}(t,\mathbf{x})}{\partial\mathbf{x}}\Bigr{|}_{{\mathbf{x}=\mathbf{x}^{*}_{i,t}}}\mathbf{x}^{*}_{i,t}),$ (20c)
$\mathbf{g}_{\mathrm{\textnormal{plan}}}(t,\mathbf{x})$ $=\begin{bmatrix}\mathbf{f}_{\mathbf{w}}\\ \mathbf{0}_{{n}_{k}}\\ ({\mathbf{f}_{\mathrm{\textnormal{plan}}}})_{({n}_{w}+1):{n}_{p}}\\ \end{bmatrix}$ (20d)
@@
Choosing a larger $\Delta t$ would induce higher mismatch between the tracking model and the planning model, thus increasing the conservativeness of the method. In scenarios with highly dynamic systems or cluttered environments, selecting a smaller $\Delta t$ reduces conservativeness, as demonstrated in Fig. 4. As will be shown in the rest of this section, PARC’s reach-avoid guarantees will still be maintained regardless of the choice of $\Delta t$, and its computation time is inversely proportional to $\Delta t$.
@@ IV-B PARC: Reach Set without Tracking Error
Without accounting for tracking error, the reach set of a planning model is the set of initial conditions $\mathbf{x}_{0}$ of trajectories that are guaranteed to reach the goal region $P_{\mathrm{\textnormal{G}}}$ at time $t_{\mathrm{\textnormal{f}}}$. Instead of naively computing the full reach set, which scales exponentially in memory, computation time, and number of set representations with the number of timesteps, PARC is planning aware: given an expert plan, we compute only a local, underapproximated BRS efficiently for the neighborhood that shares the same mode sequence as the expert plan. Fixing the mode sequence reduces the PWA system to a time-variant affine system, so the BRS is a single H-polytope using the inverse affine map of the goal (see (9)), with computation time scaling linearly with the number of timesteps and quadratically with dimensions from matrix multiplication.
@@
$\mathbf{x}(t_{\mathrm{\textnormal{f}}};\mathbf{x}_{0})\in P_{\mathrm{\textnormal{G}}}.$ (21)
Obtaining expert plans is system-dependent, but is not difficult for short-horizon and low-dimensional PWA models, because we do not require the expert plan to avoid obstacles. Thus, one option is to solve a two-point boundary value problem using nonlinear MPC [bonalli2019gusto] or mixed-integer linear program (MILP) [lofberg2004yalmip], both of which are complete. In practice, it is often sufficient to uniformly sample trajectory parameters $\mathbf{k}$ and roll out the PWA planning model, since the robot starts out near the goal. This assumption is even easier to fulfill if the planning model is a time-variant affine (which we use for a general 3D quadrotor and drift vehicle planning in Section V, VI) since plans have only one mode.
@@
$s_{t}$ $=\min\{i\ |\ \mathbf{x}(t;\mathbf{x}_{0})\in\mathcal{P}(\mathbf{A}_{i,t},\mathbf{b}_{i,t}),i=1,\dots,{n}_{\mathrm{\textnormal{PWA}},t}\},$ (22)
@@
$\begin{split}\overline{\Omega}_{t_{\mathrm{\textnormal{f}}}}&=P_{\mathrm{\textnormal{G}}},\\ \overline{\Omega}_{t}&=\mathcal{B}\!\left(\overline{\Omega}_{t+\Delta t},\mathbf{C}_{s_{t},t},\mathbf{d}_{s_{t},t}\right)\cap\mathcal{P}(\mathbf{A}_{s_{t},t},\mathbf{b}_{s_{t},t}),\end{split}$ (23)
@@
This follows directly from (LABEL:eq:time_variant_affine) and Lemma 1. ∎
@@
Traditionally, methods such as the reachableSet function in MPT3 [herceg2013multi] compute the full BRS of the PWA system instead of that of only a particular mode sequence. While computing the full BRS would return all possible reachable states and do not require knowledge of an expert plan, the computation time grows exponentially with the number of BRS steps and quickly becomes intractable due to the large number of polytopes, as shown in Table I. Instead, Proposition 10 enables a much more reasonable computation time by isolating the affine dynamics of a single mode sequence and keeping only one polytope regardless of the number of BRS steps.
[…]
@@
TABLE I: Computation time and number of polytopes for 1, 2, 3, 4, 5, and 26-step BRS for the TurtleBot example (Section III-D) using the reachableSet function in MPT3 [herceg2013multi] and Proposition 10. Since MPT3 computes the full BRS of the PWA system, the computation time quickly becomes intractable due to the large number of polytopes. In contrast, Proposition 10 maintains a reasonable computation time by keeping only one polytope at each step, characterized by the affine dynamics corresponding to the mode sequence of the expert plan.